Solve 3x+2+x+1+4=180
Answer by Jeff Gordon(65)

3x+2+x+1+4=180

combine terms

3x and x = 4x

4x+2+1+4=180

4x+7=180

subtract 7 from both sides

4x=173

x=43.25